# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ B.1.2 AIS and SMC-based score estimator
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$
@@
2Set annealing schedule $\beta_{k}\leftarrow k/n_{\mathrm{\text{steps}}}$ for $k=0,\dots,n_{\mathrm{\text{steps}}}$ ;
3 Set initial target: $\nu_{1}(u)\propto q(u\mid x_{t})^{1-\beta_{1}}\tilde{\pi}(u)^{\beta_{1}}\mathcal{N}(x_{t}\mid\alpha_{t}u,\sigma_{t}^{2}\mathbb{I})^{\beta_{1}}$ ;
4 for $m\leftarrow 1$ to $n_{\mathrm{is}}$ do
5   Sample initial particle $u^{(0,m)}\sim q(u\mid x_{t})$ ;
6   Initialize weight $w^{(m)}\leftarrow\frac{\nu_{1}(u^{(0,m)})}{q(u^{(0,m)}\mid x_{t})}$ ;
@@
8for $k\leftarrow 1$ to $n_{\mathrm{\text{steps}}}-1$ do
9   Set next target: $\nu_{k+1}(u)\propto q(u\mid x_{t})^{1-\beta_{k+1}}\tilde{\pi}(u)^{\beta_{k+1}}\mathcal{N}(x_{t}\mid\alpha_{t}u,\sigma_{t}^{2}\mathbb{I})^{\beta_{k+1}}$ ;
10 for $m\leftarrow 1$ to $n_{\mathrm{is}}$ do
11      Apply MCMC transition with kernel $\mathcal{T}$ for $m_{\mathrm{\text{steps}}}$ steps and step size $\delta_{\text{mcmc}}$ targeting $\nu_{k}(u)$ to obtain the final sample $u^{(k,m)}$ and denote the intermediate samples $\tilde{u}^{(k,m)}$;
12      Update weight: $w^{(m)}\leftarrow w^{(m)}\frac{\nu_{k+1}(u^{(k,m)})}{\nu_{k}(u^{(k,m)})}$;
@@
15Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{u^{(0,m)}\}\cup\{u^{(k,m)},\tilde{u}^{(k,m)}\}_{k=1}^{n_{\mathrm{\text{steps}}}-1}$ ;
16 Compute normalized weights: $\tilde{w}^{(m)}=\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{n_{\mathrm{is}}}w^{(m^{\prime})}}$ ;
@@
17 Compute score estimate: $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{n_{\mathrm{is}}}\tilde{w}^{(m)}\frac{\alpha_{t}u^{(n_{\mathrm{\text{steps}}}-1,m)}-x_{t}}{\sigma_{t}^{2}}$ ;
Compute marginal density estimate: $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{n_{\mathrm{is}}}\sum_{m=1}^{n_{\mathrm{is}}}w^{(m)}$;
Algorithm 3 Annealed Importance Sampler of $p(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$
Algorithm˜3 details an AIS-based approach [24] to construct the score estimate and marginal estimate.
@@
AIS introduces a sequence of intermediate unnormalized distributions, smoothly bridging a tractable initial proposal distribution $q(x_{0}\mid x_{t})$ to the posterior distribution of interest $p(x_{0}\mid x_{t})\propto\tilde{\pi}(u)\mathcal{N}(x_{t}\mid\alpha_{t}u,\sigma_{t}^{2}\mathbb{I})$. Algorithm˜3 uses a geometric annealing schedule, setting intermediate targets $\nu_{k}(x_{0})\propto q(x_{0}\mid x_{t})^{1-\beta_{k}}p(x_{0}\mid x_{t})^{\beta_{k}}$ where $\beta_{k}=k/n_{\mathrm{\text{steps}}}$ for $k=0,\cdots,n_{\mathrm{\text{steps}}}$.
@@
Algorithm˜3 begins by drawing a set of particles from the proposal distribution $q(x_{0}\mid x_{t})$. Then it gradually transitions these particles by running MCMC with a kernel invariant to the intermediate targets $\nu_{k}(x_{0})$ for $k=0,\cdots,n_{\mathrm{\text{steps}}}-1$. In this work we mainly consider Hamiltonian Monte Carlo [HMC, 45] and Langevin Dynamics [LG, 46]. At each intermediate step, the algorithm updates the importance weights of each particle based on the ratio of successive intermediate targets.
@@
The final weighted set of particles provides a consistent estimate of the desired posterior, which is utilized to construct a score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} denotes all the randomness in this procedure. Additionally, these weights provide an unbiased estimate of the unnormalized constant $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
Notably, if we resample particles based on the weights at each intermediate step, the AIS procedure effectively becomes an SMC algorithm. In that case, the incremental weight is set to $w_{k}^{(m)}\leftarrow\frac{\nu_{k+1}(u^{(k,m)})}{\nu_{k}(u^{(k,m)})}$ for $k=0,\cdots,n_{\mathrm{\text{steps}}}-1$ and $m=1,\cdots,n_{\mathrm{is}}$, and the marginal estimate becomes $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{n_{\mathrm{\text{steps}}}}\sum_{k=0}^{n_{\mathrm{\text{steps}}}-1}\frac{1}{n_{\mathrm{is}}}\sum_{m=1}^{n_{\mathrm{is}}}w_{k}^{(m)}$. The other aspect of Algorithm˜3 remains the same.
B.1.3 Proposal choices in IS/AIS-based score estimators.
@@
Algorithm˜2 and Algorithm˜3 require specifying an (initial) proposal distribution $q(x_{0}\mid x_{t})$. The choice of proposal is flexible and should ideally approximate the true posterior $p(x_{0}\mid x_{t})$; the closer it is, the more effective the resulting estimator.
In the example given in the main paper, we consider a proposal based on the backward transition kernel (i.e., the likelihood), which has also been used in prior work [8, 11, 43]. This proposal is effective when the likelihood signal is strong—i.e., when $t$ is close to 0—but becomes less efficient as $t$ increases and the signal weakens.
@@
B.2 Score estimator based on other score identities
In Eq.˜5 of the main paper, we derive the score estimator based on DSI in Eq.˜3. Alternative score estimators can be derived based on other expressions of the score.
@@ B.2 Score estimator based on other score identities
where $\{w^{(m)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{(m)}\}_{m=1}^{M}$ is a weighted system approximating the posterior $p(x_{0}\mid x_{t})$ using an IS/AIS/SMC procedure.
@@
B.3 SMC resampling and score clipping
We discuss two strategies to reduce the variance introduced by SMC resampling steps, and to address the potential high variance of score estimates, respectively.
SMC resampling strategy.
@@ B.3 SMC resampling and score clipping
We use systematic resampling, which introduces correlations among resampling indices [22, Chapter 2.2.1].
We also consider start resampling at a later stage of the sampling trajectory. At early steps, the Gaussian likelihood provides weaker conditional information, making posterior inference more challenging [8, 9]. Consequently, both the score estimates and marginal density estimates can be significantly biased or exhibit high variance, especially under limited compute budgets. As a result, the intermediate target distributions in early steps may not be sufficiently informative to enable effective resampling. To mitigate this issue, we delay resampling until a designated step $t_{\mathrm{start\text{-}resampling}}<T$, which we treat as a tunable hyperparameter. In our experiments, we found that values in the range of $[0.3T,0.8T]$ yielded improved performance.
After this point, we adopt an effective sample size (ESS)-based criterion to decide whether resampling should be performed at each step – a common strategy in SMC literature [22, Chapter 2.2.2]. The ESS at step $t$ is computed as
@@
When resampling is skipped (i.e. for $t>t_{\mathrm{start\text{-}resampling}}$ or when the normalized ESS is above $\kappa_{\mathrm{\text{ess}}}$), we must adjust the computation of the intermediate weights to ensure the unbiasedness of the log normalization constant estimate [22, Chapter 2.2.3]. For iterations where resampling is not performed, the weight update in Algorithm˜1 of Algorithm˜1 is replaced by
$w_{t}^{(i)}$ $\leftarrow\frac{\bar{w}_{t+1}^{(i)}}{1/N}\cdot\frac{\hat{p}_{t}^{(i)}p(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{p}_{t+1}^{(i)}q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$ (22)
@@
Score clipping.
To improve numerical stability, one may consider clipping the score estimates within a pre-specified range. As observed in prior work [47], score estimates produced by the TSI estimator defined in Eq.˜17 can exhibit high variance or have large magnitudes, especially during early steps. Akhound-Sadegh et al. [11] apply score norm clipping when training diffusion-based neural samplers with a loss of TSI-based score matching. Inspired by their work, we also clip the norm of the score estimates to a maximum of 20 when we use TSI estimate in the experiments. In contrast, we do not apply clipping when using the DSI-based score estimator.
B.4 Computational complexity
The total computational cost of RDSMC scales linearly with the number of diffusion discretization steps $T$, the number of particles $N$, and the cost of each score estimation step, denoted by $M$.
Specifically, for AIS-based score estimator as described in Algorithm˜3, the cost $M$ scales linearly with the number of importance samples $n_{\mathrm{is}}$, the number of AIS transition steps $n_{\mathrm{\text{steps}}}$, and the number of MCMC steps per transition $m_{\mathrm{\text{steps}}}$. The vanilla IS-based score estimator in Algorithm˜2 is a special case with $n_{\mathrm{\text{steps}}}=1$ and $m_{\mathrm{\text{steps}}}=1$.
Among these factors, the operations involving $T$, $n_{\mathrm{\text{steps}}}$, and $m_{\mathrm{\text{steps}}}$ must be processed sequentially due to their time-dependent structure. In contrast, computations over the particles $N$ and the importance samples $n_{\mathrm{is}}$ are embarrassingly parallel and can be efficiently distributed across parallel compute resources.

### v2

@@ B.1.2 AIS and SMC-based estimator
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$
@@
2Set annealing schedule $\beta_{k}\leftarrow k/n$ for $k=0,\dots,n$
3 Set initial target $\nu_{1}(x_{0})\leftarrow q(x_{0}\mid x_{t})^{1-\beta_{1}}\tilde{\pi}(x_{0})^{\beta_{1}}\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})^{\beta_{1}}$
4 for $m\leftarrow 1$ to $M$ do
5   Draw initial sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}}\sim q(\cdot\mid x_{t})$
6   Initialize weight $w^{(m)}\leftarrow\frac{\nu_{1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}})}{\nu_{0}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}})}$
@@
8for $k\leftarrow 2$ to $n$ do
9   Set the current target $\nu_{k}(x_{0})\propto q(x_{0}\mid x_{t})^{1-\beta_{k}}\tilde{\pi}(x_{0})^{\beta_{k}}\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})^{\beta_{k}}$
10 for $m\leftarrow 1$ to $M$ do
11      Update sample $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}$} using MCMC transition targeting $\nu_{k}$ starting from $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k-1,m)}$}
12      Update weight $w^{(m)}\leftarrow w^{(m)}\frac{\nu_{k}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}})}{\nu_{k-1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}})}$
@@
15Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}}\}_{m=1}^{M}\}_{k=1}^{n}$
16 Compute normalized weights $\tilde{w}^{(m)}\leftarrow\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}$
@@
17 Compute score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{M}\tilde{w}^{(m)}\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}}-x_{t}}{\sigma_{t}^{2}}$
Compute marginal density estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}w^{(m)}$
Algorithm 3 Annealed Importance Sampling (AIS)-based Score and Marginal Estimator
@@
Algorithm˜3 presents an AIS-based approach [neal2001annealed] to construct the score and marginal estimates. AIS introduces a sequence of intermediate unnormalized distributions, smoothly bridging a tractable initial proposal distribution $q(x_{0}\mid x_{t})$ to the denoising posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. Under a geometric annealing schedule, intermediate targets are $\nu_{k}(x_{0})\propto q(x_{0}\mid x_{t})^{1-\beta_{k}}\pi(x_{0}\mid x_{t})^{\beta_{k}}$ where $\beta_{k}=k/n$ for $k=0,\cdots,n$.
@@
Algorithm˜3 begins by drawing a set of particles from the proposal distribution $q(x_{0}\mid x_{t})$. Then it gradually transitions these particles by running MCMC with a kernel invariant to the intermediate targets $\nu_{k+1}(x_{0})$ for $k=0,\cdots,n-1$. In this work we mainly consider Metropolis-Hasting adjusted Langevin dynamics [MALA, roberts1996exponential] and Hamiltonian Monte Carlo [HMC, neal2012mcmc]. At each intermediate step, the algorithm updates the importance weights of each particle based on the ratio of successive intermediate targets.
@@
The final weighted set of particles provides a consistent estimate of the desired posterior, which is utilized to construct a score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} denotes all the randomness in this procedure. Additionally, these weights provide an unbiased estimate of marginal density $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, i.e. $\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]=Z\pi(x_{t})$, where $q(x_{t}\mid x_{t})$ is the sampling distribution of $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} given $x_{t}$.
SMC adaption of Algorithm˜3.
We can adapt Algorithm˜3 to an SMC-based procedure by incorporating the resampling mechanism. In this case, the weight update in Algorithm˜3 is modified to the incremental weight computation:
$w^{(k,m)}\leftarrow\frac{\nu_{k+1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}{\nu_{k}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}$
and the initial weight in Algorithm˜3 is denoted by $w^{(1,m)}$, for $m=1,\cdots,M$. With this adjustment, we resample particles $\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$ in the beginning of each iteration $k$ based on the previous weights $\{w^{(k-1,m)}\}_{m=1}^{M}$ for $k=2,\cdots,n$. The marginal estimate in Algorithm˜3 becomes
$\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{n}\sum_{k=1}^{n}\frac{1}{M}\sum_{m=1}^{M}w^{(k,m)}.$
The other aspect of Algorithm˜3 remains the same.
B.1.3 Rejection sampling (RS)-based estimator
[…]
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
1
[…]
3 while $m\leq M$ do
4   Sample proposal ${\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\sim q(\cdot\mid x_{t})$
5   Compute acceptance ratio $r\leftarrow\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}v_{i}})\,\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}},\sigma_{t}^{2}I)}{C\,q({\color[rgb]{0.3125,0.46875,0.62890625}v_{i}})}$
6   Draw ${\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\sim\mathrm{Uniform}(0,1)$
7 if $v<r$ then
8      Accept sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\leftarrow{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}$
9    $m\leftarrow m+1$
10
11 $i\leftarrow i+1$
12
[…]
14 Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}},{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\}_{i=1}^{M_{\textrm{total}}}$
15 Compute score estimate: $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}}$
16 Compute marginal estimate: $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow C\frac{M}{M_{\textrm{total}}}$
Algorithm 4 Rejection Sampling (RS)–based score and marginal estimator
In Algorithm˜4, we present a rejection sampling (RS)-based approach, building on prior work of he2024zeroth which propose an RS-based score estimator and develop theoretical guarantees for the resulting diffusion MC sampler. We extend their approach to additionally provide an unbiased marginal density estimate alongside the original score estimate. In principle, this procedure can be used within RDSMC (Algorithm˜1) as a replacement for the AIS scheme (Algorithm˜3). Future work may explore this direction particularly in low-dimensional settings, as he2024zeroth demonstrate advantages over other score estimators in a non-SMC context.
Notably, rejection sampling requires access to a constant $C$ that upper bounds the ratio of the unnormalized target to the proposal. he2024zeroth use Newton’s method to find a proxy of $C$.
B.1.4 Proposal choices in IS/AIS/SMC/RS-based score estimators.
@@ 11 $i\leftarrow i+1$
Algorithms˜2, 3 and 4 require specifying an (initial) proposal distribution $q(x_{0}\mid x_{t})$. The choice of this proposal is flexible and should ideally approximate the true posterior $\pi(x_{0}\mid x_{t})$; the closer it is, the more effective the resulting estimator.
In Algorithms˜2, 3 and 4, the default initial proposal is obtained by reversing the forward noising likelihood, suggested by prior work [huang2023reverse, akhound2024iterated, he2024training]. This proposal is effective when the likelihood signal is strong—i.e., when $t$ is close to 0—but becomes less efficient as $t$ increases and the signal weakens.
@@
B.1.5 Score estimator based on other score identities
In Eq.˜5 of the main paper, we derive the score estimator based on the DSI in Eq.˜3. Alternative score estimators can be derived based on other score identities.
@@ B.1.5 Score estimator based on other score identities
where $\{w^{(m)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{(m)}\}_{m=1}^{M}$ is a weighted system approximating the posterior $\pi(x_{0}\mid x_{t})$ using an IS/AIS/SMC procedure.
@@
B.2 Computational complexity
The total computational cost of RDSMC scales linearly with the number of diffusion discretization steps $T$, the number of particles $N$, and the cost of each score estimation step.
Specifically, for AIS-based score estimator as described in Algorithm˜3, the score estimation cost scales linearly with the number of importance samples $M$, the number of AIS transition steps $n$, and the number of MCMC steps per transition $l$. The vanilla IS-based score estimator in Algorithm˜2 is a special case with $n=1$ and $l=1$. Therefore the cost of score estimation per SMC particle per time step $\mathcal{O}(Mnl)$. Consequently, the overall complexity for RDSMC is $\mathcal{O}(NT\cdot Mnl)$.
Among these factors, the operations involving discretization steps $T$, AIS transition steps $n$, and MCMC steps $l$ must be processed sequentially due to their time-dependent structure. In contrast, computations over the particles $N$ and the importance samples $M$ are embarrassingly parallel and can be efficiently distributed across parallel compute resources.
B.3 Implementation guidelines
We discuss practical implementation guidelines for RDSMC, including the resampling strategy in §˜B.3.1, score clipping in §˜B.3.2 and hyperparameter choices in §˜B.3.3.
B.3.1 SMC resampling strategy
@@
First, we use systematic resampling, which introduces correlations among resampling indices [naesseth2019elements, Chapter 2.2.1].
We also consider start resampling at a later stage of the sampling trajectory. At early steps, the Gaussian likelihood provides weaker conditional information, making posterior inference more challenging [huang2023reverse, grenioux2024stochastic]. Consequently, both the score estimates and marginal density estimates can be significantly biased or exhibit high variance, especially under limited compute budgets. As a result, the intermediate target distributions in early steps may not be sufficiently informative to enable effective resampling. To mitigate this issue, we delay resampling until a designated step $t_{\mathrm{start\text{-}resampling}}<T$, which we treat as a tunable hyperparameter.
For timestep $t<t_{\mathrm{start\text{-}resampling}}$, we adopt an effective sample size (ESS)-based criterion to decide whether resampling should be performed at each step – a common strategy in SMC literature [naesseth2019elements, Chapter 2.2.2]. The ESS at step $t$ is computed as
@@
When resampling is skipped (i.e. for $t>t_{\mathrm{start\text{-}resampling}}$ or when the normalized ESS is above $\kappa_{\mathrm{\text{ess}}}$), we must adjust the computation of the intermediate weights to ensure the unbiasedness of the log normalization constant estimate [naesseth2019elements, Chapter 2.2.3]. For iterations where resampling is not performed, the weight update in Algorithm˜1 of Algorithm˜1 is replaced by
$w_{t}^{(i)}$ $\leftarrow\frac{\bar{w}_{t+1}^{(i)}}{1/N}\cdot\frac{\hat{\pi}_{t}^{(i)}p(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$ (22)
@@
B.3.2 Score clipping
To improve numerical stability, one may consider clipping the score estimates within a pre-specified range. Note that score clipping does not affect our theoretical guarantees provided that regularity conditions hold, as the score estimate is only part of the proposal mechanism.
As observed in prior work [noble2024learned], score estimates produced by the TSI estimator defined in Eq.˜17 can exhibit high variance or have large magnitudes, especially during early steps. akhound2024iterated apply score norm clipping when training diffusion samplers where the loss function is based on the TSI. Inspired by their work, we also clip the norm of the score estimates to a maximum of 20 when we use the TSI-based score estimates. In contrast, we do not apply clipping when using the DSI-based score estimator.
B.3.3 Hyperparameter choices
The main hyperparameters of RDSMC are:
Diffusion-related hyperparameters.
  • Diffusion parameterization: we use the VP diffusion schedule described in Eq.˜16 in this work.
  • Time discretization strategy: we use a uniform time discretization grid with $T=100$ steps, though more informative strategies can be incorporated—such as the SNR-adapted scheme proposed by grenioux2024stochastic.
MC score estimation-related hyperparameters.
  • Score estimator type: we use AIS-based score estimator (Algorithm˜3), with vanilla IS (Algorithm˜2) as a special case. We mainly leverage the DSI (Eq.˜5) to estimate the score in most experiments, while we also use the TSI (Eq.˜19) in Bayesian logistic regression tasks.
  • MCMC transition kernel: MALA or HMC.
  • $M$: Number of importance samples used in Monte Carlo score estimation
  • $n$: Number of intermediate distributions, i.e. number of AIS steps; for IS $n=1$
  • $\delta_{\text{mcmc}}$: Step size used in the MCMC transition kernel (not used in IS). We manually tune this parameter in the main experiments while we use adaptive step size in ablation studies; see discussion below.
  • $l$ : Number of MCMC steps per AIS transition (not used in IS). We use $l=1$ in all experiments.
SMC-related hyperparameters.
  • $t_{\mathrm{start\text{-}resampling}}$: Starting time for resampling within the SMC trajectory. No resampling is performed for $t/T>t_{\mathrm{start\text{-}resampling}}$.
  • $\kappa_{\mathrm{\text{ess}}}$: Normalized effective sample size threshold triggering resampling. When $\kappa_{\mathrm{\text{ess}}}=0$, no resampling is applied throughout, which corresponds to RDSMC (IS) if a final-step weighting is applied and RDSMC (Proposal) otherwise. We set $\kappa_{\mathrm{\text{ess}}}=0.3$ in all experiments for RDSMC.
In the main experiments, we choose score estimator and MCMC kernel types by a coarse grid search, and tune $M,n,\delta_{\text{mcmc}}$ and $t_{\mathrm{start\text{-}resampling}}$ in a range of values, using validation metrics. The hyperparameter values are summarized Table˜3; see details in §˜E.1.2.
In the ablation experiments in §˜E.5, we fix most hyperparameters to reduce the degrees of freedom. We use $T=100,M=100,n=50,m=1,\kappa_{\mathrm{\text{ess}}}=0.3$, and AIS-based score estimator leveraging the DSI, LG as the MCMC kernel. We also use an adaptive MCMC step size $\delta_{\text{mcmc}}$, as in grenioux2024stochastic. When the acceptance rate is too high ($76\%$) we increase the step size by $3\%$, when it is $74\%$, we decrease step size by $3\%$, and otherwise we keep the current step size. Only an initial step size value needs to be specified and then it is adpatively adjusted during the sampling process. The only hyperparameter in RDSMC that requires tuning is $t_{\mathrm{start\text{-}resampling}}$ , which we vary in $\{0,0.1,\cdots,0.9,1.0\}$. We observe consistent performance and therefore in practice we would recommend users to start with something similar to this configuration.